If a 0.16 kg ball has a momentum of 5.36 kg٠m/s as it is thrown from a machine to the right. What is its velocity? 


1
Expert's answer
2021-03-29T09:02:25-0400
"p=mv\\v=\frac{p}{m}=\frac{5.36}{0.16}=33.5\frac{m}{s}"